A jar of quarters and dimes contains $15.25. their are 103 coins in all. how many of each are their?
Answer by checkley77(12569)

Q+D=103 OR Q=103-D
.25Q+.10D=15.25
.25(103-D)+.10D=15.25
25.75-.25D+.10D=15.25
-.15D=15.25-25.75
-.15D=10.50
D=10.50/.15
D=70 NUMBER OF DIMES.
103-70=33 NUMBER OF QUARTERS.
PROOF:
.25*33+.10*70=15.25
8.25+7=15.25
15.15=15.25